# A quantum computing concept for 1-D elastic wave simulation

### v1

@@ 4 Implementation and experiments on quantum simulators and computers
The transformations discussed in the previous sections are pre-processing steps executed on a classical computer. In the following, we present the implementation of the transformed elastic wave equation on both an actual QC and an ideal (error-free) quantum simulator. A quantum simulator is a software running on classical computers that emulates the behaviour of actual QCs. We create a quantum circuit that defines all operations that must be physically conducted on qubits during the circuit’s runtime. All of these operations, except for the measurements of the qubit states, are necessarily unitary. Hence, the quantum circuit itself can be represented by a single unitary matrix.
The quantum circuit that solves the elastic wave equation through quantum time evolution is composed of 3 sub-circuits; (1) one that prepares the initial condition |$\ket{\Phi(0)}$, (2) one that performs the quantum simulation through implementing the Hamiltonian $\mathbf{H}$, and (3) one that utilises the evolved state |$\ket{\Phi(t)}$. This utilisation can, for example, be a measurement of the qubits yielding binary results. In this section, we employ the open-source software development kit Qiskit by IBM Research (Aleksandrowicz et al., 2019) to create and run the quantum circuit on quantum simulators and actual QCs. The representation of the initial conditions |$\ket{\Phi(0)}$ on a QC requires normalisation,
|$\ket{\Phi_{\mathrm{norm}}(0)}=\frac{\ket{\Phi(0)}}{||\ket{\Phi(0)}||}\,,$ (26)
because quantum states must have a unit norm. This normalisation has to be undone at the end of the simulation via
|$\ket{\Phi(t)}=\ket{\Phi_{\mathrm{norm}}(t)}||\ket{\Phi(0)}||\,.$ | | . (27)
@@
|$\ket{\Phi_{\mathrm{norm}}(0)}=\mathbf{V}\ket{00..00}\,.$ (28)
This is done by the Qiskit method qc.prepare_state(phi0_norm). Constructing and applying the quantum circuit corresponding to $\mathbf{V}$ can be challenging (e.g., Kak, 1999). However, in the context of seismic applications and using the above-derived transformation $\mathbf{T}$, the initial state |$\ket{\Phi_{\mathrm{norm}}(0)}$ is sparse and can be initialised efficiently as discussed in appendix B. Next, the quantum wave simulation circuit is realised by appending the time-evolution operator
@@
|$\ket{\Phi_{\mathrm{norm}}(t)}=\mathbf{U}(t)\ket{\Phi_{\mathrm{norm}}(0)}.$ (30)
We currently perform the time evolution, i.e., the action of $\mathbf{U}$, using Qiskit class PauliEvolutionGate. It implements the scaling and squaring method (Al-Mohy and Higham, 2010). We have chosen this method because it is implementable on today’s freely accessible QCs, such as IBM Brisbane, with a moderate circuit overhead for small-scale problems. For large-scale problems, which will only become tractable on QCs with higher qubit counts and better error correction, this approach becomes inapplicable due to the exponential scaling; for $n$ qubits, $\mathbf{U}$ is a $2^{n}\times 2^{n}$ dense matrix that needs to be stored and applied to |$\ket{\Phi_{\mathrm{norm}}(0)}$. As already explained in section 2.3, the action of $\mathbf{U}$ must be carried out only approximately to obtain a scalable approach. Appendix B discusses scaling if the best available approaches are applied to our settings.
Finally, the resulting quantum state |$\ket{\Phi(t)}$ is utilised further within the QC or is read out using quantum state tomography (Smolin et al., 2012). Here, we opt for the latter to compare the QC simulation with the solution obtained on a classical computer. However, this becomes prohibitively expensive for large problems. Notably, in situations where the state vectors are strictly real, |$\ket{\Phi(t)}$ forms a $2^{n}$-dimensional distribution that needs to be sampled. Additionally, to get a sample, a quantum measurement needs to be performed, which collapses the state (Nielsen and Chuang, 2010), therefore requiring a new full run for each sample. Reading out the final state is then more computationally demanding than the HS itself. We discuss how to circumvent this problem in the context of imaging in section 5. In summary, the quantum time evolution circuit proceeds as follows:
@@
    State tomography via repeated measurement and evaluation.
@@
The quantum state tomography output, |$\ket{\Phi_{\mathrm{norm}}(t)}$, represents the quantum state after the simulation. To retrieve the classical wave field, we reverse the normalisation, apply the inverse transform $\mathbf{T}^{-1}$ and subsequently apply $\mathbf{M}^{-1/2}$.
For demonstration, we simulate a 7 grid-point problem, which we found to be the maximum number of grid points realisable on IBM Brisbane. The setup and all results of the numerical experiment are summarised in Fig. 1. We use Dirichlet boundary conditions, a uniform space discretisation $dx$ of 1 m, and monotonically increasing density and elastic modulus distributions.
As the numerical ground truth, we use an ordinary differential equation (ODE) solver based on an explicit Runge-Kutta method (e.g., Dormand and Prince, 1980; Shampine, 1986) to directly integrate (33). Since the quantum formalism is exactly isomorphic to the classical one, the HS should exactly reproduce this solution, including finite-difference errors such as grid dispersion caused by the small number of grid points per wavelength. To investigate the correctness of our quantum implementation, we use an ideal QC simulator, i.e., a quantum simulator that does not imitate quantum noise and decoherence effects of the current-generation QCs but appears as perfectly noise-free. The numerical ground truth agrees with the solution from the simulator. A relative error of 0.01 %, which is not visible in the plot, is attributed to the quantum tomography, as discussed above.
Finally, we solve the problem on the real QC IBM Brisbane. To perform the quantum state tomography extracting the quantum simulation results, we run the whole circuit $10^{3}\times 2^{n}$ times to reduce the relative error contribution of the quantum state tomography to $0.03$. We observe qualitative agreement between the numerical ground truth and the simulation on the QC but substantial errors of up to 60 % in the norm of their difference. These can be attributed to the high decoherence and noise rates of real QCs, as opposed to the ideal QC simulator.
@@
When a quantum algorithm is executed, a sequence of quantum gates is applied to each qubit as discussed in section 2. The number of gates applied to the qubits is known as the circuit depth. Each gate application has a certain success rate, the probability that the gate operation is executed correctly. IBM Brisbane’s gates have an error per layered gate for a 100-qubit chain (EPLG) of 1.9 %, equaling an approximate average 5-qubit layer success rate of 99.9 % (Gambetta (2020); Collins and Easterly (2021)). For the 7 grid point problem, our implementation yields an average circuit depth of 870 quantum gates with 5 qubits, which results roughly in a probability of $0.999^{870}\approx 0.42$ for a successful execution of the whole circuit. This indeed reflects approximately the error level of 60 % in our simulation. Increasing the number of grid points will increase the circuit depth and hence the error, currently limiting the problem size to 7 grid points.
On today’s QCs, the error characteristics vary substantially per hardware and problem. Quantum processing units (QPUs) have different interaction topographies, implemented gate sets, and gate error rates of individual qubits and require continuous re-calibration. Hence, it is expected that different error characteristics would appear if we had chosen a different QC, different qubits within the same QC, or even just a different runtime. Improving the reliability of quantum chips, in combination with more advanced error mitigation methods, is a vivid research field. It is, therefore, expected that substantial improvements will be achieved in the coming years and decades.
@@
Figure 1: Comparison of elastic wave field simulations using a classical ODE solver (black), an ideal quantum simulator (red), and the quantum computer IBM Brisbane (blue). (a) Heterogeneous distributions of density and elastic modulus. (b-e) wave field at selected times.
@@ 5 Discussion
In the following paragraphs, we provide additional details concerning potential computational speed-ups, the utilisation of the final state, a continuous version of the above-describe transformations in the context of alternative numerical methods, and future extensions of the 1-D elastic concept.
@@ 5.1 Exponential quantum speedup
The first promise of quantum computing concerns the storage size. For classical computational methods, the number of bits required to store the entire wave field scales linearly with the number of grid points. In contrast, for quantum computational methods like the one discussed here, the number of required qubits scales only logarithmically because, reciprocally, the quantum state space grows exponentially with the number of qubits. For instance, assume a discretised wave field with two real-valued parameters per grid point, such as displacement and velocity, and Earth volume $1.083\times 10^{12}\ \textrm{km}^{3}$. Representing a full wave field in $\textrm{mm}^{3}$ resolution requires merely 101 qubits.
The second promise is runtime speed-up. Building on work by Babbush et al. (2023), we demonstrate in appendix B, that the quantum algorithm presented in sections 2.34 has a runtime complexity that is polynomial in $n$ and poly-logarithmic in $N$, provided that the best currently available approaches are used for the individual simulation steps. In contrast, any classical algorithm necessarily has at least exponential complexity in $n$ and polynomial in $N$ (Babbush et al., 2023). In the foreseeable future, the specific $N$ for which the quantum approach becomes actually faster might be above the limit given by the available hardware. However, the promise is highly enticing in the long term.
5.2 Utilising the final state
As discussed in section 4, reading out the full state with quantum state tomography scales polynomially in $N$ and exponentially in $n$. Hence, the read-out would nullify the exponential speed-up. Babbush et al. (2023) show that it is possible to read out certain properties, such as kinetic energy or potential energy of a subdomain of the model, in polynomial time, suggesting that the exponential speed-up can be preserved in such cases. A similar logic will be required to preserve the exponential speed-up for seismological purposes, e.g., by computing a misfit function on the fly instead of reading out the complete time evolution of the wave field. This topic will be the subject of a follow-up publication.
@@ 5.3 Continuous transformation and the use of other numerical schemes
In appendix A, we show that there exists such a transformation which is invertible. The resulting Hamiltonian is given by
@@
It is now possible to discretise the continuous Schrödinger equation directly by discretising $\mu^{1/2}\partial_{x}\rho^{-1/2}$ with any numerical method, provided that the boundary conditions can be satisfied.
5.4 Extensions
The work presented here is merely a first step towards QC wave simulations in geophysical applications. Generalising to higher space dimensions will require a similar procedure as above to map the augmented wave operator $\mathcal{Q}$ or its discretised version $\mathbf{Q}$ to the Hamiltonian operator, ensuring self-adjointness under the quantum inner product. A bigger challenge may be the incorporation of visco-elastic attenuation because HS inherently conserves energy.

### v2

@@ 4 Implementation and experiments on quantum simulators and computers
The transformations discussed in the previous sections are pre-processing steps executed on a classical computer. In the following, we present the implementation of the transformed elastic wave equation on both an actual QC and an ideal (error-free) quantum simulator. A quantum simulator is a software running on classical computers that emulates the behavior of actual QCs. We create a quantum circuit that defines all operations that must be physically conducted on qubits during the circuit’s runtime. All of these operations, except for the measurements of the qubit states, are necessarily unitary. Hence, the quantum circuit itself can be represented by a single unitary matrix.
The quantum circuit that solves the elastic wave equation through quantum time evolution is composed of 3 sub-circuits; (1) one that prepares the initial condition |$\ket{\phi(0)}$, (2) one that performs the quantum simulation through implementing the Hamiltonian $\mathbf{H}$, and (3) one that utilises the evolved state |$\ket{\phi(t)}$. This utilization can, for example, be a measurement of the qubits yielding binary results. In this section, we employ the open-source software development kit Qiskit by IBM Research (Aleksandrowicz et al., 2019) to create and run the quantum circuit on quantum simulators and actual QCs. The representation of the initial conditions |$\ket{\phi(0)}$ on a QC requires normalisation,
|$\ket{\psi(0)}=\frac{\ket{\phi(0)}}{||\ket{\phi(0)}||}\,,$ (26)
because quantum states must have a unit norm. This normalization has to be undone at the end of the simulation via
|$\ket{\phi(t)}=\ket{\psi(t)}||\ket{\phi(0)}||\,.$ | | . (27)
@@
|$\ket{\psi(0)}=\mathbf{V}\ket{00..00}\,.$ (28)
This is done by the Qiskit method qc.prepare_state(psi0). Constructing and applying the quantum circuit corresponding to $\mathbf{V}$ can be challenging (e.g., Kak, 1999). However, in the context of seismic applications and using the above-derived transformation $\mathbf{T}$, the initial state |$\ket{\psi(0)}$ is sparse and can be initialised efficiently as discussed in appendix B. Next, the quantum wave simulation circuit is realized by appending the time-evolution operator
@@
|$\ket{\psi(t)}=\mathbf{U}(t)\ket{\psi(0)}.$ (30)
We currently perform the time evolution, i.e., the action of $\mathbf{U}$, using the Qiskit class PauliEvolutionGate. It implements the scaling and squaring method (Al-Mohy and Higham, 2010). The unitary time evolution for a particular evolution time, $\mathbf{U}(t)$, is constructed classically and decomposed into quantum gates. These gates are implemented as part of the overall quantum circuit. Therefore, the time evolution is equivalent to a single matrix-vector product between the time evolution operator $\mathbf{U}(t)$ and the initial state |$\ket{\psi(0)}$. We have chosen this method because it is implementable on today’s freely accessible QCs, such as IBM Brisbane, with a moderate circuit overhead for small-scale problems. For large-scale problems, which will only become tractable on QCs with higher qubit counts and better error correction, this approach becomes inapplicable due to the exponential scaling; for $n$ qubits, $\mathbf{U}$ is a $2^{n}\times 2^{n}$ dense matrix that needs to be stored and applied to |$\ket{\psi(0)}$. As already explained in section 2.3, the action of $\mathbf{U}$ must be carried out only approximately to obtain a scalable approach. Appendix B discusses scaling if the best available approaches are applied to our settings.
Finally, the resulting quantum state |$\ket{\psi(t)}$ is utilized further within the QC or is read out using quantum state tomography (QST) (Smolin et al., 2012). Here, we opt for the latter to compare the QC simulation with the solution obtained on a classical computer. As discussed shortly, this approach is not practical for larger problems but is used here to validate the theory and implementation.
The squared entries $\{|\psi_{i}(t)|^{2}\}_{i=1}^{2^{n}}$ define a probability mass function which can be sampled by running and measuring the circuit multiple times. QST allows us to go beyond sampling the probability mass function but instead infer the actual amplitudes $\{\psi_{i}(t)\}_{i=1}^{2^{n}}$. It includes two steps: Firstly, estimating the expectation value of $3^{n}$ Pauli observables, which in our case reduces to $2^{n}$ because |$\ket{\psi(t)}$ is purely real. Secondly, after all measurements are acquired, they are used in a least-squares inversion process to fit the amplitudes of |$\ket{\psi(t)}$ by determining the density matrix |$\ket{\psi(t)}\bra{\psi(t)}$ | and its eigenvectors and eigenvalues (Smolin et al., 2012). Finally, it should be noted that the error in the estimated amplitudes $\{\psi^{est.}_{i}(t)\}_{i=1}^{2^{n}}$, given by $\epsilon=|\psi_{i}(t)-\psi^{est.}_{i}(t)|$, decreases up to prefactors with $1/\sqrt{S}$, where $S$ is the number of samples.
QST is prohibitively expensive for large problems, due to the exponential, here $3^{n}$ $(2^{n})$ number of observables that need to be measured for a general complex (real) state. Additionally, to obtain a single sample with respect to an observable, a quantum measurement needs to be performed, which collapses the state (Nielsen and Chuang, 2010), therefore requiring a new HS run for each sample. Reading out the final state therefore requires exponentially many HS passes, nullifying any speedup from HS. We discuss this problem and perspectives to resolve it in the context of imaging in section 5.
In summary, the quantum time evolution circuit proceeds as follows:
@@
    Quantum state tomography (QST) via repeated HS runs and measurements.
Details on the implemented quantum circuit using Qiskit can be found in Appendix D.
@@
The quantum state tomography output, |$\ket{\psi(t)}$, represents the quantum state after the simulation. To retrieve the classical wave field $[\mathbf{u},\mathbf{v}]$, we reverse the normalization, apply the inverse transform $\mathbf{T}^{-1}$, and subsequently apply $\mathbf{M}^{-1/2}$.
We use a QC simulator which is run on a classical computer to investigate the correctness of our algorithm. There are two types of simulators: one where artificial noise is added to simulate the real behavior of QCs and one which is noise-free. We opt for the latter to test the theory. Therefore, any divergence from the ground truth is attributed to finite sampling noise in the QST procedure. Only in the limit of infinitely many samples, QST reproduces the actual quantum state. As the numerical ground truth, we use an ordinary differential equation (ODE) solver based on an explicit Runge-Kutta method (e.g., Dormand and Prince, 1980; Shampine, 1986) to directly integrate equation (34). We use Dirichlet boundary conditions, a uniform space discretization $dx$ of 1 m, and monotonically increasing density and elastic modulus distributions. Since the quantum formalism is isomorphic to the discretized classical one, the HS should reproduce this solution, including finite-difference errors such as e.g. numerical dispersion. Figure 1 displays the simulation results of the classical ground truth (black circles) and of quantum simulators using 20 (orange) and 1000 (red) samples for the QST, respectively. We use a relative $L_{2}$ error (RL2) of the wave field amplitudes $u$ as a comparison metric calculated as
$E(t)=\frac{\sqrt{\sum_{i}^{N}(u_{i}(t)-u^{g}_{i}(t))^{2}}}{\sqrt{\sum_{i}^{N}u^{g}_{i}(t)^{2}}},$ (31)
with $u^{g}$ the ground truth solution obtained with the ODE solver. For the time steps shown in Figure 1, the RL2 error is roughly 20 % and 4 % when $20$ and $1000$ samples are used, respectively. Including even more samples would reduce the error further with a scaling of $1/\sqrt{S}$, as discussed above. Because we use a matrix exponential scheme that evolves the state with a single matrix application (see Section 4), the error does not accumulate over time but exclusively results from the QST and is determined by the statistics of the probability mass function of the state at the particular time point. However, for more elaborate time evolution algorithms that can be implemented on future QCs, such as the one discussed in Appendix B, the error will increase over time.
Finally, we solve the problem on the real QC IBM Brisbane, requiring us to scale the problem down to $7$ grid points, as discussed shortly. The results are displayed in Figure 2. We observe qualitative agreement between the numerical ground truth and the simulation on IBM Brisbane, however with RL2 errors of up to 60 %. These errors can be attributed to the high decoherence and noise rates on real QCs, as quantum states are extremely fragile and hard to isolate from the environment. We measure each quantum circuit with 1000 samples, reducing the RL2 contribution of the QST to approximately 3 %, which is far below the hardware noise level that dominates the error. Again, the error does not accumulate over time, as discussed above, but rather depends on the quantum circuit for a particular time step and its hardware implementation.
@@
When a quantum algorithm is executed, a sequence of quantum gates is applied to each qubit as discussed in section 2. The number of gates applied to the qubits is known as the circuit depth. Each gate application has a certain success rate, the probability that the gate operation is executed correctly. IBM Brisbane’s gates have an error per layered gate for a 100-qubit chain (EPLG) of 1.9 %, equaling an approximate average 5-qubit layer success rate of 99.9 % (Gambetta (2020); Collins and Easterly (2021)). For the 7 grid point problem, our implementation yields an average circuit depth of 870 quantum gates with 5 qubits, which results roughly in a probability of $0.999^{870}\approx 0.42$ for a successful execution of the whole circuit. This indeed reflects approximately the error level of 60 % in our simulation. Increasing the number of grid points will increase the circuit depth and hence the error beyond the noise level, currently limiting the problem size to 7 grid points. This is not a limit imposed by the presented algorithm, but a limitation of currently available hardware.
The error characteristics of QCs vary substantially per hardware and problem. Different quantum processing units have different interaction topographies, implemented gate sets, and gate error rates of individual qubits and require continuous, often daily, re-calibration which changes the error characteristics. Because the problem specifics such as the number of grid points, evolution time, material parameters, initial conditions, or boundary conditions alter the quantum circuit, they influence the error. However, since currently the error mostly depends on the hardware which is calibrated and updated rapidly, a detailed problem-specific error analysis is not meaningful at this stage.
While currently HS on QCs is still in the proof-of-principle state, improving the reliability of quantum chips, in combination with more advanced error mitigation methods, is a vivid research field. It is expected that substantial improvements will be achieved in the coming years and decades (Gambetta (2020); Kim et al. (2023)). Additionally, it has recently been demonstrated that HS is tractable on noisy QCs (Kikuchi et al. (2023)), which may open avenues for applications of HS in the short term.
@@
Figure 1: Comparison of elastic wave field simulations using a classical ODE solver (black) and two ideal quantum simulators with 20 samples (orange) and 1000 samples (red) used in the respective QST for a 128 grid-point problem. (a) Heterogeneous distributions of density and elastic modulus. (b-e) wave field at selected times.
Figure 2: Comparison of elastic wave field simulations using a classical ODE solver (black), an ideal quantum simulator (red), and the quantum computer IBM Brisbane (blue) with 1000 samples used in the QST for a 7 grid-point problem. (a) Heterogeneous distributions of density and elastic modulus. (b-e) wave field at selected times.
@@ 5 Discussion
In the following paragraphs, we provide additional details concerning potential computational speedups, the utilization of the final state, a continuous version of the above-described transformations in the context of alternative numerical methods, and future extensions of the 1-D elastic concept.
@@ 5.1 Exponential quantum speedup
The first promise of quantum computing concerns the storage size. For classical computational methods, the number of bits required to store the entire wave field scales linearly with the number of grid points. In contrast, for quantum computational methods like the one discussed here, the number of required qubits scales only logarithmically because, reciprocally, the quantum state space grows exponentially with the number of qubits. For instance, assume a discretized wave field with two real-valued parameters per grid point, such as displacement and velocity, and Earth volume $1.083\times 10^{12}\ \textrm{km}^{3}$. Representing a full wave field in $\textrm{mm}^{3}$ resolution requires merely 101 qubits.
The second promise is runtime speedup. As discussed in Section 4, the time evolution method implemented in our QC experiment has been chosen because it has a small circuit overhead, allowing for implementation on current hardware. This formulation does not provide any speedup over classical computers as it scales exponentially in the number of qubits. However, as detailed in Appendix B, the best quantum algorithms for HS have a runtime complexity that is polynomial in $n$ and poly-logarithmic in $N$ (Low and Chuang (2019); Babbush et al. (2023)). We analyze those algorithms only theoretically as they are not implementable on current hardware due to the large circuit overhead.
In contrast to the best HS algorithms, any classical algorithm necessarily has at least exponential complexity in $n$ and polynomial in $N$ (Babbush et al., 2023). In the foreseeable future, the specific $N$ for which the quantum approach becomes faster might be above the limit given by the available hardware. However, the promise is highly enticing in the long term.
5.2 The read-out problem
As discussed in section 4, reading out the full state with QST scales polynomially in $N$ and exponentially in $n$. Hence, the read-out would nullify the exponential speedup. Babbush et al. (2023) show that it is possible to read out certain properties, such as kinetic energy or potential energy of a subdomain of the model, in polynomial time, suggesting that the exponential speedup can be preserved in such cases. A similar logic will be required to preserve the exponential speedup for seismological purposes, e.g., by computing a misfit function on the fly instead of reading out the complete time evolution of the wave field. This topic will be the subject of a follow-up publication.
@@ 5.3 Continuous transformation and the use of other numerical schemes
In appendix A, we show that there exists such a transformation that is invertible. The resulting Hamiltonian is given by
@@
It is now possible to discretize the continuous Schrödinger equation directly by discretizing $\mu^{1/2}\partial_{x}\rho^{-1/2}$ with any numerical method, provided that the boundary conditions can be satisfied.
5.4 Extensions and limitations
The work presented here is merely a first step toward QC wave simulations in geophysical applications. Generalizing to higher space dimensions will require a similar procedure as above to map the augmented wave operator $\mathcal{Q}$ or its discretized version $\mathbf{Q}$ to the Hamiltonian operator, ensuring self-adjointness under the quantum inner product. A bigger challenge may be the incorporation of visco-elastic attenuation because HS is inherently unitary and hence can only simulate systems that conserve energy. Another open problem is the incorporation of sources, which will be the topic of a follow-up study.
5.5 Relation to other quantum algorithms with potential for seismological research
In this article, the QC was used as a semi-analog simulator through HS. There exist other families of algorithms that have the potential to be harnessed in seismology. Among the most prominent are variational quantum algorithms (Cerezo et al. (2021)) which use a hybrid quantum-classical optimization procedure to find eigenvalues and eigenvectors (Abrams and Lloyd (1999); Kirby and Love (2021)) or solve linear systems (Pellow-Jarman et al. (2021)). The former may have applications in solving the wave equation in the frequency domain. The latter has been used to solve linear partial differential equations, including the wave equation (Trahan et al. (2023)), by successively solving the linear system of equations that arises in implicit time stepping schemes. While their computational benefits in comparison to classical solutions and HS still need to be worked out, they bring the advantage of being executable on today’s and near-term quantum chips of the NISQ era (Cerezo et al. (2021)).
Again looking ahead into the potential future of error-corrected quantum chips, Harrow et al. (2009) have proposed a quantum algorithm, known as the HHL algorithm, that is capable of solving sparse linear systems with an exponential speedup over classical computers. Such algorithms can eventually be harnessed in a discrete-time integration scheme similar to Trahan et al. (2023) or in linearized inverse problems (Wiebe et al. (2012); Schuld et al. (2016)). Furthermore, despite concerns over vanishing gradients issues (McClean et al. (2018); Wang et al. (2021)), quantum machine learning (Biamonte et al. (2017); Melnikov et al. (2023)) is an emerging field with potential application in seismological research.
Finally, there are quantum annealers (QAs) or adiabatic quantum computers (Albash and Lidar (2018a)), which differ from gate-based QCs. QAs are capable of solving quadratic optimization problems and their application in geoscience has been investigated in a handful of studies (O’Malley (2018); Golden and O’Malley (2021); Souza et al. (2022); Dukalski et al. (2023)). It should be noted that there is controversy about whether or not QA can achieve the same speedup as gate-based quantum computation in real-world contexts (Van Dam et al. (2001); Aharonov et al. (2008); Albash and Lidar (2018b); Villanueva et al. (2023)).